OpenStudy (anonymous):

evaluate the expression for x-15 , y=4 2x + 18y

OpenStudy (across):

I've never been this confused before. Could you please repeat the question?

OpenStudy (anonymous):

Can someone help solve this equation? 2x+18y=

OpenStudy (across):

2x+18y=...?

OpenStudy (anonymous):

yeah that lol

OpenStudy (across):

What's supposed to be on the right hand side?

OpenStudy (karatechopper):

plug 4 in for y and mult and then add

OpenStudy (anonymous):

thats how is showed sorry

OpenStudy (karatechopper):

yo answer would b -3 and x+ 36

OpenStudy (anonymous):

evaluate the expression for x= 19, y=6 x+9/y+8

OpenStudy (anonymous):

x-15 , y=4 2x + 18y yeah you take and plug in the values of the given numbers for the letters that are given so like x=-15 you take and put -15 where x is and same thing for y.

OpenStudy (karatechopper):

2

OpenStudy (anonymous):

evaluate the expression for a=15,b=25,c=26,d=1 a+(b+c) x d